OpenStudy (anonymous):

pls explain for me why (3/2pi)=sin2x => x= (1/2)arcsin(3/2pi)

OpenStudy (anonymous):

\[\frac{ 3 }{ 2\pi }=\sin2x => x=\frac{ 1 }{ 2 }\arcsin(\frac{ 3 }{ 2\pi })\]

OpenStudy (anonymous):

cos(3y) dy = -sin(2x) dx Integrating both sides, we have INT cos(3y) dy = INT -sin(2x) dx (1/3)sin(3y) = (1/2)cos(2x) + C Given the initial condition y(pi/2) = pi/3, solve for C. (1/3)sin(3*pi/3) = (1/2)cos(2*pi/2) + C 0 = (1/2)*(-1) + C C = 1/2

OpenStudy (anonymous):

@dangbeau there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay\[\frac{ 3 }{ 2\pi } =\sin(2x)\] taking sine inverse we get \[\sin^{-1}\frac{ 3 }{ 2\pi } = 2x\] we want x right so subjecting it we get\[x = \frac{ 1 }{2}\sin^{-1} (\frac{ 3 }{ 2\pi })\]

OpenStudy (anonymous):

IS it clear to u now ?

OpenStudy (anonymous):

i got it, thank you ^^